OpenStudy (anonymous):

37.051 is 55.3% of what number

OpenStudy (anonymous):

\[x \times 0.553 =37.051\rightarrow x=37.051/0.553=67\]

OpenStudy (anonymous):

or do someting like 37.051 ----- 55.3 x ----- 100\[x=37.051\times100\div55.3=67\]

OpenStudy (anonymous):

why do you have to bump the decimal over like that

OpenStudy (anonymous):

55.3%=55.3/100=0.553